# Binaural Target Speaker Extraction using HRTFs

### v3

@@ Abstract
In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers. To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker. Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets and languages. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals, and compare it to a Real-Imaginary (RI)-based neural network, demonstrating the advantages of the former. We first evaluate the method in an anechoic, noise-free scenario, achieving excellent extraction performance while preserving the binaural cues of the target signal. We then extend the evaluation to reverberant conditions. Our method proves robust, maintaining speech clarity and source directionality while simultaneously reducing reverberation. A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods demonstrates that our approach attains performance on par with competing techniques in terms of noise reduction and perceptual quality, while offering a clear advantage in preserving binaural cues. Demo-page: https://bi-ctse-hrtf.github.io
Index Terms—  Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
1 Introduction
@@
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims to selectively extract the desired speaker from multiple competing speakers in a binaural setting. In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids employ either a single microphone or multiple microphones per earpiece and can be broadly categorized into two groups: 1) Bilateral systems, where each device operates independently and provides a single-channel output per earpiece; 2) Shared systems, where all microphones across both earpieces are jointly utilized, yielding a dual-channel output.
@@
Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears. In bilateral systems, where each device processes its own channel independently, it is particularly challenging to preserve these binaural cues. In contrast, Multi input Multi output (MIMO) solutions jointly process both channels, yielding more natural and spatially accurate outputs. For example, [4] presents a Binaural Linearly Constrained Minimum Variance (BLCMV) beamformer that separates sources while maintaining binaural cues. More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
@@
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge of the target speaker(s) to isolate their voice. This prior knowledge—commonly referred to as a “clue” [6]—can take various forms. A widely used clue is a pre-recorded speech sample, often referred to as an enrollment utterance, which provides speaker-specific features [7, 8, 9, 10]. Clues may also come from other modalities, such as visual data. For example, [11] uses a face image to extract the corresponding speech, while [12] combines audio and visual clues, with the audio clue being an enrollment utterance and the visual clue a video of the speaker speaking within the mixture.
@@
While examining prior works, such as [13], we observed a particular challenge in separating two voices with highly similar characteristics, for example, speakers of the same gender. Consequently, we opted for an approach that does not depend on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA and extracts spatial features from it. These spatially guided methods consistently outperform single-microphone solutions.
@@
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction cue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
2 Problem Formulation
@@
We address the problem of multiple concurrent speakers, captured by two microphones, each mounted on one of the listener’s ears in a noiseless environment. For clarity, we focus on the simpler case of two concurrent speakers, as formulated below.
@@
Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation). Similarly, the interfering speaker’s DOA is ($\theta_{i}$$\phi_{i}$). Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively. The binaural signal can be written as
@@
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech signals of the desired speaker and the interfering speaker, respectively. Let $\gamma\in\{d,i\}$ denote the speaker index, where $d$ refers to the desired source and $i$ to the interferer. The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all reflections. The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizes the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$.
@@
According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$. Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
@@
Given the HRTF corresponding to the desired source’s DOA $(\theta_{d},\phi_{d})$, namely the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. The extracted signal is denoted by $\hat{\tilde{\bm{s}}}_{d}$. This binaural output is required to: 1) preserve the spatial cues of the desired source, and 2) dereverberate the output by retaining only the direct-path HRTF, thereby also dereverberating the desired speech signal.
3 Proposed Method
@@
We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
3.1 Architecture
@@
The proposed model architecture is based on the approach presented in [14], modified to fully handle dual-channel, complex-valued inputs, rather than the real and imaginary components independently. The proposed model employs a U-Net architecture, incorporating the HRTF at the bottleneck. The individual modules comprising the architecture are described in detail below.
@@
Encoder-Decoder Blocks of the U-net: The encoder comprises five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation function:
$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\left(\mathcal{R}(z)\right)+\mathrm{j}\cdot\text{ReLU}\left(\mathcal{I}(z)\right).$ (2)
@@
Here $\text{ReLU}(x)=\max(0,x)$ and $\mathcal{R}(z),\mathcal{I}(z)$ denote the real and imaginary parts of z, respectively. Additionally, $\mathrm{j}$ represents the imaginary unit. Similarly, the decoder comprises five up-convolution blocks, each employing transposed convolutional layers. Skip connections are utilized between corresponding encoder and decoder blocks of the U-net structure to improve reconstruction accuracy. Additionally, the decoder includes a final convolutional layer with no activation function, enabling the representation of both positive and negative values at the output. This design choice is essential since the STFT representation inherently includes both positive and negative real and imaginary components.
Bottelneck Processing: The extraction clue is applied to the encoded input features at the bottleneck, with self-attention layers employed both before and after its application. However, unlike traditional methods, we implement these self-attention layers directly on complex-valued tensors by replacing the softmax activation function with the complex-valued $\mathbb{C}\text{ReLU}$.
To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor undergoes four consecutive self-attention layers and a fully connected layer, preparing it for the decoding stage. A block diagram of the proposed architecture appears in Fig. 1.
@@
Fig. 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated desired signal.
3.2 Loss function
We initially trained the model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
@@ 3.2 Loss function
where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all samples of the target and estimated signals, respectively. However, relying solely on the SI-SDR loss proved challenging in preserving accurate spectral details, resulting in noticeable musical noise artifacts in the extracted signals. Musical noise refers to tonal distortions that occur due to inaccuracies or rapid variations in the estimated signal’s frequency-domain representation [17]. Such artifacts are commonly observed in signal enhancement tasks when spectral continuity and consistency are not adequately enforced. To mitigate these artifacts, we introduced an additional frequency-domain loss, the $L1$- Mean Absolute Error (MAE) loss:
@@
Here, $c=\{l,r\}$ denotes the channels, left or right. To improve data utilization, both speakers were extracted from each mixed signal in each training step by alternating the HRTF clues and averaging the SI-SDR losses. The overall training loss is therefore defined in (5), as a weighted combination of the SI-SDR and $L1$-MAE losses:
@@
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality. The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
4 Experimental Study
@@ 4.1 Data Generation
where $L$ is the reflection order. The direct path corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$.
The RIEC HRTF Dataset [19, 20]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA. We employed the SofaMyRoom framework [23] to simulate both anechoic and reverberant conditions, following the formulation in (6). This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time. It also supports zero-order simulations for anechoic scenarios. The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room. The overall simulation setup is illustrated in Fig. 2.
Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [24]. To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
@@
Fig. 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$. Images source: freepik.com
4.2 Algorithm Setup

### v4

# Yoav Ellinson    Sharon Gannot

@@ Abstract
In this work, we address the problem of binaural target-speaker extraction in the presence of multiple simultaneous talkers. We propose a novel approach that leverages the individual listener’s Head-Related Transfer Function (HRTF) to isolate the target speaker. The proposed method is speaker-independent, as it does not rely on speaker embeddings. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals, and compare it to a Real-Imaginary (RI)-based neural network, demonstrating the advantages of the former. We first evaluate the method in an anechoic, noise-free scenario, achieving excellent extraction performance while preserving the binaural cues of the target signal. We then extend the evaluation to reverberant conditions. Our method proves robust, maintaining speech clarity and source directionality while simultaneously reducing reverberation. A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods shows that the proposed approach achieves performance comparable to state-of-the-art techniques in terms of noise reduction and perceptual quality, while providing a clear advantage in preserving binaural cues. Demo-page: https://bi-ctse-hrtf.github.io
I Introduction
@@
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments with multiple speakers. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims to selectively extract the desired speaker from multiple competing speakers in a binaural setting. In this context, preserving the binaural cues is critical: hearing aids and their associated algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids employ either a single microphone or multiple microphones per earpiece and can be broadly categorized into two groups: 1) Bilateral systems, where each device operates independently and provides a single-channel output per earpiece; 2) Shared systems, where all microphones across both earpieces are jointly utilized, yielding a dual-channel output.
@@
Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears. In bilateral systems, in which each device processes its own channel independently, it is particularly challenging to preserve binaural cues. In contrast, Multi input Multi output (MIMO) solutions jointly process both channels, yielding more natural and spatially accurate outputs. For example, [4] presents a Binaural Linearly Constrained Minimum Variance (BLCMV) beamformer that separates sources while maintaining binaural cues. More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
@@
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge on the target speaker(s) to isolate their voice. This prior knowledge—commonly referred to as a “clue” [6]—can take various forms. A widely used clue is a pre-recorded speech sample, often referred to as an enrollment utterance, which provides speaker-specific features [7, 8, 9, 10]. Clues may also come from other modalities, such as visual data. For example, [11] uses a face image to extract the corresponding speech, while [12] combines audio and visual clues, with the audio clue being an enrollment utterance and the visual clue a video of the speaker speaking within the mixture.
@@
While examining prior works, such as [13], we observed a particular challenge in separating two voices with highly similar characteristics, for example, speakers of the same gender. Consequently, we opted for an approach that does not rely on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and more effectively isolate the desired speaker. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15, 16] rely on the DOA of the desired speaker and use it as a spatial clue. These spatially guided methods consistently outperform single-microphone solutions.
@@
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the individual listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction clue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct path’s HRTF, thereby dereverberating the signal and hence further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
II Problem Formulation
@@
We address the problem of multiple concurrent speakers captured by two microphones mounted on the listener’s ears, assuming a noiseless environment and a fixed source–listener radius. For clarity, we focus on the case of two concurrent speakers, as formulated below.
@@
Denote the desired speaker’s DOA by $(\theta_{d},\phi_{d})$ and the interferer’s by $(\theta_{i},\phi_{i})$. Let $\bm{x}_{b}(n,k)$ denote the STFT of the binaural signal, with $n\in[0,N-1]$ and $k\in[0,K-1]$ the time-frame and frequency-bin indices, respectively. The binaural signal is therefore formulated as:
@@
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech signals of the desired speaker and the interfering speaker, respectively. Let $\gamma\in\{d,i\}$ denote the speaker index, where $d$ refers to the desired source and $i$ to the interferer. The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all acoustic reflections in the enclosure, with the HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizeing the listener’s response to a wave arriving from DOA $(\theta,\phi)$.
@@
According to the precedence effect [17], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$. Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
@@
Given the individual HRTF of the hearing-aid wearer corresponding to the desired source DOA $(\theta_{d},\phi_{d})$, denoted by the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. The extracted signal is denoted by $\hat{\tilde{\bm{s}}}_{d}$. It is required that the binaural output satisfy the following conditions: 1) preservation of the spatial cues of the desired source, and 2) dereverberation of the desired speech output by retaining only the direct-path HRTF. In this work, we focus on individualized HRTF database, namely the set of HRTFs of the hearing aid wearer across all azimuth and elevation directions, which should be given to the algorithm. HRTF ersonalization using anthropometric measurements [18].
III Proposed Method
@@
We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue.
III-A Architecture
@@
The proposed model architecture is built upon the approach presented in [14], with modifications to handle dual-channel, complex-valued inputs rather than treating the real and imaginary components independently. The proposed model employs a U-Net architecture, incorporating the HRTF at the bottleneck. The individual modules comprising the architecture are described in detail below and depicted in Fig. 1.
@@
Encoder–Decoder Blocks of the U-Net: The encoder consists of five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation defined as
$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\!\left(\mathcal{R}(z)\right)+\mathrm{j}\,\text{ReLU}\!\left(\mathcal{I}(z)\right),$ (2)
@@
where $\text{ReLU}(x)=\max(0,x)$, $\mathcal{R}(z)$ and $\mathcal{I}(z)$ denote the real and imaginary parts of $z$, and $\mathrm{j}$ is the imaginary unit. The decoder mirrors the encoder with five up-convolution blocks based on transposed convolutions. Skip connections are employed between corresponding encoder and decoder blocks to improve reconstruction accuracy. The decoder further includes a final convolutional layer without an activation function, allowing the output to represent both positive and negative real and imaginary values, which is required for the complex STFT representation.
Bottelneck: The extraction clue is applied at the bottleneck of the encoded input features, with self-attention layers used both before and after its application. Unlike conventional approaches, these self-attention layers operate directly on complex-valued tensors by replacing the softmax activation with the complex-valued $\mathbb{C}\text{ReLU}$. To capture cross-channel (left–right) information in the HRTF, a self-attention mechanism is employed, enabling the channels to attend to one another via learned weights. The resulting matrix is flattened and projected via a fully-connected layer to match the dimensionality of the encoded mixture, yielding the HRTF embedding. Similarly, self-attention is applied to the encoded mixed signal. The replicated HRTF embedding is then combined with each time frame of the encoded mixture through element-wise multiplication, thereby imposing the extraction clue. The resulting tensor is processed by four additional self-attention layers and a fully connected layer before decoding.
@@
Figure 1: A block diagram of the proposed method, where $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the HRTF of the desired speaker’s DOA, $\bm{x}_{b}$ represents the mixed binaural signal, and $\hat{\tilde{\bm{s}}}_{d}$ represents the estimated desired signal, both in the STFT domain.
III-B Loss function
The model was initially trained using only the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss (3).
@@ III-B Loss function
where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all time-domain samples of the target and estimated signals, respectively. However, relying solely on the SI-SDR loss proved challenging in preserving accurate spectral details, resulting in noticeable musical noise artifacts in the extracted signals. Musical noise refers to tonal distortions that occur due to inaccuracies or rapid variations in the estimated signal’s frequency-domain representation [19]. Such artifacts are commonly observed in signal enhancement tasks when spectral continuity and consistency are not adequately enforced. To mitigate these artifacts, we introduced an additional frequency-domain loss, the $L1$- Mean Absolute Error (MAE) loss:
@@
Here, $c=\{l,r\}$ denotes the output channels, left or right. To improve data utilization, both speakers were extracted from each mixed signal in each training step by alternating the HRTF clues and averaging the SI-SDR losses. The overall training loss is therefore defined in (5), as a weighted combination of the SI-SDR and $L1$-MAE losses:
@@
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, in which we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality. The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
IV Experimental Study
@@ IV-A Data Generation
where $L$ is the reflection order. The direct path ($\ell=0$) corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$, and the gain $a_{\ell}$ and the time delay $\tau_{\ell}$, obtained from the image method, are applied by scaling and time-shifting each HRTF component.
The RIEC HRTF Dataset [21, 22]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [23, 24] facilitates easy retrieval of the appropriate HRTF for any desired DOA. We employed the SofaMyRoom framework [25] to simulate both anechoic and reverberant conditions, following the formulation in (6). This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time. It also supports zero-order simulations for anechoic scenarios. The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and the listener position is drawn uniformly within the room. The overall simulation setup is illustrated in Fig. 2.
Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [26].
@@
Figure 2: A simulation illustration: two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$. Images source: freepik.com
IV-B Algorithm Setup